OpenStudy (anonymous):

elimination method a+7b=13 a+2b=-2

OpenStudy (anonymous):

7b-2b = 13-(-2) 5b = 15 b = 3

OpenStudy (anonymous):

a + 7(3) = 13 a + 21 = 13 a = 13-21 a=-8

OpenStudy (anonymous):

So the one solution is? 3,-8?

OpenStudy (anonymous):

\[a+7b=13[1]\] \[a+2b=-2[2]\] \[[1]-[2]\] \[(a+7b)-(a+2b)=(13)-(-2)\] \[a+7b-a-2b=13--2\] \[5b=13+2\] \[5b=15\] \[b=3\] Find a by subbing b as 3 into one of the equations given to you above.

OpenStudy (anonymous):

Make sure you don't just seek for the answers. Actually look at how people solve the questions. Look at their methods and their approach to the question. You will better understand how to solve these if there comes a time for an exam.

OpenStudy (anonymous):

Definitely. I am trying to better understand the entire concept

OpenStudy (anonymous):

since the point is finding a and b then the answer is an ordered pair starting with a correct?

OpenStudy (anonymous):

Could you rephrase that last part please? ordered pair starting with a?

OpenStudy (anonymous):

a is 3 so it is 3,-8

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no it's not 3.

OpenStudy (anonymous):

You mixed it up.

OpenStudy (anonymous):

I meant te othe way around sorry, -8,3)

OpenStudy (anonymous):

3=b